# Kineo: Calibration-Free Metric Motion Capture From Sparse RGB Cameras

### v1

@@ 2.2.2 Camera Extrinsics
2.4 Output Representation and Performance Considerations
@@
A major source of complexity in this optimization arises from the type of 3D representation used. Parametric models [27, 12, 33] are widely adopted due to their compactness and strong structural priors. These models encode pose and shape in low-dimensional spaces, enabling them to represent plausible pose and shape configurations even under occlusions or depth ambiguities. Such priors are particularly valuable in single-view settings, where geometric constraints are crucial to resolve ambiguity. However, in multiview scenarios, where ambiguities are naturally reduced through additional viewpoints, these structural constraints become less critical. The computational cost of parametric models remains high due to the multiple transformation stages required (e.g., pose decoding from latent space[33], blendshape evaluation, linear blend skinning), as well as the iterative fitting needed to align model projections with 2D image evidence for each video frame. Moreover, these models are typically constrained to specific subject categories, such as humans, and are subject to biases like mean-shape bias [41, 39].
By contrast, non-parametric models, including keypoints, voxels, point clouds, etc., offer a more direct and computationally efficient alternative but do not provide inherent constraints. Keypoint-based representations, in particular, are widely used in multiview and optimization-based pipelines due to their lightweight nature. While they lack explicit structural constraints, soft priors can be introduced to guide the optimization toward plausible solutions [5]. In cases where such priors are omitted, we refer to the models as free-form non-parametric representations. A hybrid strategy has also emerged, wherein non-parametric predictions are converted to parametric models parameters for compactness and use in downstream tasks. In particular, Sárándi et al. [39] propose a fast method to fit SMPL meshes to dense non-parametric 3D keypoints.
Hybrid pipelines that combine learning-based regressors with subsequent optimization have shown strong performance and enhanced robustness in challenging scenarios. A representative example is HSfM, a calibration-free method that achieves good accuracy by leveraging multiple specialized models to jointly reconstruct cameras, scenes and human subjects. Specifically, HSfM employs SAM2[37] for segmentation, HMR2.0[9] for pose regression, DUSt3R[55] to estimate the scene and cameras, ViTPose[58] for keypoint detection, and SMPL[27] for body modeling. These components provide complementary information that, when refined through a global optimization stage, lead to SOTA quality reconstructions. However, this architecture also increases computational complexity and results in longer runtimes due to the integration of several heavyweight models.
@@
3 Method overview
Figure 1: Overview of the Kineo pipeline for markerless motion capture from uncalibrated and unsynchronized multi-camera videos. Starting from raw video inputs, the system first performs audio-based temporal synchronization to align unsynchronized streams on a shared timeline. Next, automatic camera calibration estimates extrinsic and intrinsic parameters, including Brown–Conrady lens distortion, via a graph-based optimization over 2D keypoint correspondences selected by a confidence-driven sampling strategy. Using the recovered cameras, 3D keypoints and scene point maps are reconstructed, and each triangulated keypoint is assigned a pairwise reprojection confidence score to quantify reconstruction quality. Finally, metric-scale recovery is achieved either through a human body prior using the SMPL model or a monocular metric depth estimator for subject-agnostic scaling. The modular design enables robust and scalable reconstruction of both human and non-human subjects across long, multi-view sequences.
@@ 3 Method overview
To obtain metric-scale reconstructions, we propose two strategies. The first leverages a human body prior, using the SMPL model when people are present in the scene, similar in concept to HSfM [29]. The second is subject-agnostic, employing a monocular metric depth estimator to recover scale even in scenes without humans (see Sec. 7).
4 Temporal Synchronization
@@
Rather than relying on raw spectrogram cross-correlation, which is computationally demanding for long recordings [11, 43], we use compact audio descriptors to improve efficiency. Specifically, we represent each audio track using Mel-Frequency Cepstral Coefficients (MFCCs), a perceptually motivated and widely adopted feature in audio analysis. MFCCs capture the spectral envelope of the signal and are robust to noise. The reduced set of features allows for efficient and accurate cross-correlation, making them a suitable choice for estimating temporal lag in multi-camera settings.
The accuracy of this method depends on two main factors: the temporal resolution of the MFCC representation and the physical distance between the sound source and each camera. The former is determined by the hop length $H$ used during the Short-Time Fourier Transform (STFT), and the latter by the time delay resulting from different distances to the sound source. Following prior observations [11], we formalize these two sources of error as $\Delta t_{1}$ and $\Delta t_{2}$, respectively. We derive conditions under which subframe synchronization is achievable. Let $f_{s}$ be the audio sample rate, $f_{r}$ the video frame rate, $H$ the hop length, $d_{1}$ and $d_{2}$ the distance from each camera to the sound source respectively, and $c\penalty 10000\ \approx 343\,\text{m/s}$ the speed of sound in air. Then:
$\Delta t_{1}=\frac{H}{2f_{s}}\penalty 10000\ \quad\penalty 10000\ \text{and}\penalty 10000\ \quad\penalty 10000\ \Delta t_{2}=\penalty 10000\ \frac{|d_{1}-d_{2}|}{c}$
@@
$\Delta t_{1}+\Delta t_{2}<\penalty 10000\ \frac{1}{f_{r}}$
@@
$H<\penalty 10000\ \frac{2f_{s}}{f_{r}}-\frac{2f_{s}|d_{1}-d_{2}|}{c}\quad\text{and}\quad|d_{1}-d_{2}|<c(\frac{1}{f_{r}}-\frac{H}{2f_{s}})$
@@
In typical multi-view camera arrangements (e.g., circular rigs), microphone distances to the sound source vary by only a few meters. For instance, with $|d_{1}-d_{2}|<5$ m, $f_{s}=48$ kHz, and $f_{r}=60$ Hz, the hop length $H$ must be below 200 samples to ensure subframe accuracy. Inversely, with $H=1$, the method remains accurate for inter-microphone distances below approximately 5.7 m. In case finer alignment is required, image-based synchronization techniques [52, 44] can be used to refine an initial audio-based estimate. In all experiments presented in this work, the camera setups fall within our constraints, and audio-based synchronization is sufficient to achieve subframe accuracy. As a result, we did not incorporate image-based refinement in our pipeline.
5 Graph-based Automatic Camera Calibration
@@
This graph formulation enables global reasoning about pairwise relationships and allows us to select an optimal subset of calibrations through a minimum spanning tree (MST). This contrasts with prior incremental approaches that add cameras greedily to a global frame, at the risk of yielding a suboptimal global configuration [56, 34]. By composing relative poses along the MST, we recover global camera extrinsics efficiently with a guarantee of minimizing the total calibration error (Section 5.5). The recovered cameras are expressed in arbitrary units, and metric scaling to real-world dimensions is applied in Sec. 7.
5.1 Intrinsics Initialization
@@
To construct the camera graph, we need to recover the relative poses between camera pairs, which requires decomposing the essential matrices into their corresponding rotations and translation directions. Computing the essential matrix, however, depends on knowledge of the camera intrinsics $K$. In our scenario, where cameras are assumed to be static and the subject is moving, Kruppa’s formulation [23, 10] cannot be applied. One alternative is to compute fundamental matrices, which do not require intrinsics, and then optimize the intrinsics for each camera by minimizing the Sampson distance over all edges. However, full-graph optimization is computationally expensive, growing exponentially with the number of cameras. To maintain a linear cost with respect to the number of cameras, we instead use MoGe [54] to obtain an initial estimate of each camera’s intrinsics. This model is later reused to predict scene point maps. These intrinsics are later refined during bundle adjustment (Sec. 5.6), to which we incorporate lens distortion coefficients that follow the Brown-Conrady model, a step omitted in prior works.
5.2 Efficient Pairwise Calibration
@@
We propose a novel approach for pairwise camera calibration that is both robust and scalable to long sequences due to its fixed computational cost. Our method builds upon the work of Puwein et al. [34] by introducing confidence-driven spatio-temporal subsampling of keypoint correspondences, which enhances robustness while maintaining efficiency. Additionally, we leverage a graph-based formulation to compute relative scales reliably, incorporating the quality of each pairwise calibration to produce geometrically consistent relative extrinsics.
@@
The edges of the graph represent pairwise camera calibrations, encoded by an essential matrix $E_{ij}$ that captures the relative rotation $R_{ij}$ and translation direction $\hat{\mathbf{t}}_{ij}$ between cameras $C_{i}$ and $C_{j}$. Classical approaches estimate this matrix using the Structure-from-Motion (SfM) five-point algorithm [31] within a RANSAC loop applied either to all raw keypoint correspondences [56] or to a subset of correspondences [34]. Both strategies can be unstable when low-quality correspondences are present, and processing all correspondences becomes computationally prohibitive for long sequences.
@@
Learning-based methods such as DUSt3r [55, 29] predict relative poses in a single-frame multiview setting and can be more robust when high-confidence keypoints are unavailable. However, they do not exploit temporal consistency, which can limit accuracy, and their reliance on an additional neural network increases computational overhead; for example, DUSt3r requires approximately 40 ms for inference on a single image pair using an H100 GPU. In contrast, our method is more computationally frugal because it fully exploits the same 2D keypoints that are already required for triangulating the final skeleton output. By selectively subsampling these keypoints based on their detection confidence, we improve the stability and accuracy of previous approaches while preserving the efficiency of classical SfM and ensuring scalability to long sequences.
5.2.1 Confidence-Aware Spatio-Temporal Subsampling
@@
Figure 2: Example confidence-driven keypoint subsampling for a pair of views from Human3.6M. Colors indicate the pair confidence score $w_{kl}$, solid lines represent selected correspondences, and dashed lines represent correspondences that were not selected. The sampling favors keypoints for which the 2D keypoint detector is highly confident about their location (e.g., unoccluded and well-localized keypoints).
@@
An effective spatio-temporal subsampling strategy for keypoint correspondences is important for three reasons: (1) it improves the robustness of essential matrix estimation by filtering out low-quality correspondences and by leveraging information across multiple frames, (2) it reduces computational complexity by limiting the number of correspondences, which is particularly important for scalability on long sequences, and (3) it preserves sufficient geometric diversity to ensure accurate essential matrix estimation. Traditional RANSAC samples correspondences uniformly, ignoring the varying reliability of detections. As a result, if low-confidence correspondences constitute a majority, they can disproportionately influence hypothesis generation, leading to inaccurate relative pose estimates and requiring more iterations to converge on a valid model. Puwein et al. [34] addressed this by sampling correspondences with higher probabilities for temporally distant correspondences and for those involving different keypoints. However, their strategy does not leverage keypoint confidence scores, which can lead to the inclusion of unreliable correspondences.
@@
Building on their method, we improve the spatio-temporal subsampling strategy by incorporating the 2D keypoints detector confidence scores $w_{i,k}\penalty 10000\ \in[0,1]$, with $i$ the view index and $k$ the keypoints index, into the selection process. Each point correspondence between camera $C_{i}$ and $C_{j}$ is denoted as $(\mathbf{x}_{i,k},\mathbf{x}_{j,k})$. Its score is computed as the geometric mean of the detector confidences:
$w_{ij,k}=\sqrt{w_{i,k}\penalty 10000\ \cdot w_{j,k}}$ (1)
@@
5.3 Graph Edge Reliability Estimation
@@
$\hat{\mathbf{x}}_{i,k}=K_{i}^{-1}\mathbf{x}_{i,k}\penalty 10000\ \quad\penalty 10000\ \text{and}\penalty 10000\ \quad\penalty 10000\ \hat{\mathbf{x}}_{j,k}=K_{j}^{-1}\mathbf{x}_{j,k}$
@@
$s_{ij}=\frac{1}{M}\sum_{m=1}^{M}\mathcal{S}(E_{ij},\hat{\mathbf{x}}_{i,m},\hat{\mathbf{x}}_{j,m})$ (2)
@@
5.4 Solving relative scales
@@
$R_{j}=R_{i}R_{ij}\penalty 10000\ \quad\penalty 10000\ \text{and}\penalty 10000\ \quad\penalty 10000\ \mathbf{t_{j}}=R_{i}\cdot(\lambda_{ij}\cdot\penalty 10000\ \mathbf{\hat{t_{ij}}})+\mathbf{t_{i}}$
In previous work, finding this relative scale $\lambda_{ij}$ is addressed by aligning 3D structures in a common space. Xu et al.[56] rescale and align the triangulated bones from each camera pair into a shared coordinate system. In their method, the quality of the pairwise triangulation directly affects the result when computing the relative scales. Methods like DUSt3r[55] incorporate the scaling factors into a global alignment loop. The predicted pairwise point maps are then scaled and aligned into a globally consistent 3D reconstruction.
Instead, we solve for the relative scales $\lambda_{ij}$ by exploiting loop closure constraints [2]. The key idea is that if we follow the relative transformations along a closed cycle of cameras, for example moving from $C_{i}$ to $C_{j}$ to $C_{k}$ and back to $C_{i}$, we should return to the starting point. This concept is illustrated in Fig. 3. Formally, each triplet of cameras yields one linear constraint of the form:
$\penalty 10000\ \lambda_{ij}R_{ki}R_{jk}\penalty 10000\ \mathbf{\hat{t_{ij}}}+\penalty 10000\ \lambda_{jk}R_{ki}\penalty 10000\ \mathbf{\hat{t_{jk}}}+\penalty 10000\ \lambda_{ki}\penalty 10000\ \mathbf{\hat{t_{ki}}}=\penalty 10000\ \mathbf{0}$ (3)
We collect all $L$ loop constraints into a homogeneous linear system $A\boldsymbol{\lambda}=\mathbf{0}$, where $A\in\mathbb{R}^{3L\times M}$ contains the coefficients from the constraints, and $\boldsymbol{\lambda}\penalty 10000\ \in\penalty 10000\ \mathbb{R}^{M}$ represents the unknown relative scale factors.
@@
$WA\boldsymbol{\lambda}=\mathbf{0}$ (5)
@@
$\min_{\boldsymbol{\lambda}}\|W^{\frac{1}{2}}A\boldsymbol{\lambda}\|_{2}^{2}+\mu\|\boldsymbol{\lambda}-\mathbf{1}\|_{2}^{2}$ (6)
@@
Figure 3: Example of camera loop closure with incorrect (left) and correct (right) relative scale factors. When the relative scales $\lambda_{ij}$ are incorrect, traversing the cycle $C_{i}\!\to\!C_{j}\!\to\!C_{k}\!\to\!C_{i}^{\prime}$ does not return to the original camera position, resulting in $C_{i}\neq C_{i}^{\prime}$. In contrast, when the correct scales are used, the composition of the relative transformations yields the identity, and the loop closes consistently with $C_{i}=C_{i}^{\prime}$.
5.5 Optimal Absolute Extrinsics

### v2

@@ 2.2.2. Camera Extrinsics
2.4. Output Representation and Performance Considerations
@@
A major source of complexity in this optimization arises from the type of 3D representation used. Parametric models (Loper et al., 2015; Hewitt et al., 2023; Pavlakos et al., 2019) are widely adopted due to their compactness and strong structural priors. These models encode pose and shape in low-dimensional spaces, enabling them to represent plausible pose and shape configurations even under occlusions or depth ambiguities. Such priors are particularly valuable in single-view settings, where geometric constraints are crucial to resolve ambiguity. However, in multiview scenarios, where ambiguities are naturally reduced through additional viewpoints, these structural constraints become less critical. The computational cost of parametric models remains high due to the multiple transformation stages required (e.g., pose decoding from latent space(Pavlakos et al., 2019), blendshape evaluation, linear blend skinning), as well as the iterative fitting needed to align model projections with 2D image evidence for each video frame. Moreover, these models are typically constrained to specific subject categories, such as humans, and are subject to biases like mean-shape bias (Sengupta et al., 2020; Sárándi and Pons-Moll, 2025).
By contrast, non-parametric models, including keypoints, voxels, point clouds, etc., offer a more direct and computationally efficient alternative but do not provide inherent constraints. Keypoint-based representations, in particular, are widely used in multiview and optimization-based pipelines due to their lightweight nature. While they lack explicit structural constraints, soft priors can be introduced to guide the optimization toward plausible solutions (Bragagnolo et al., 2025). In cases where such priors are omitted, we refer to the models as free-form non-parametric representations. A hybrid strategy has also emerged, wherein non-parametric predictions are converted to parametric models parameters for compactness and use in downstream tasks. In particular, Sárándi et al. (Sárándi and Pons-Moll, 2025) propose a fast method to fit SMPL meshes to dense non-parametric 3D keypoints.
Hybrid pipelines that combine learning-based regressors with subsequent optimization have shown strong performance and enhanced robustness in challenging scenarios. A representative example is HSfM, a calibration-free method that achieves good accuracy by leveraging multiple specialized models to jointly reconstruct cameras, scenes and human subjects. Specifically, HSfM employs SAM2(Ravi et al., 2024) for segmentation, HMR2.0(Goel et al., 2023) for pose regression, DUSt3R(Wang et al., 2024) to estimate the scene and cameras, ViTPose(Xu et al., 2022) for keypoint detection, and SMPL(Loper et al., 2015) for body modeling. These components provide complementary information that, when refined through a global optimization stage, lead to SOTA quality reconstructions. However, this architecture also increases computational complexity and results in longer runtimes due to the integration of several heavyweight models.
@@
3. Method overview
Figure 1. Overview of the Kineo pipeline for markerless motion capture from uncalibrated and unsynchronized multi-camera videos. Starting from raw video inputs, the system first performs audio-based temporal synchronization to align unsynchronized streams on a shared timeline. Next, automatic camera calibration estimates extrinsic and intrinsic parameters, including Brown–Conrady lens distortion, via a graph-based optimization over 2D keypoint correspondences selected by a confidence-driven sampling strategy. Using the recovered cameras, 3D keypoints and scene point maps are reconstructed, and each triangulated keypoint is assigned a pairwise reprojection confidence score to quantify reconstruction quality. Finally, metric-scale recovery is achieved either through a human body prior using the SMPL model or a monocular metric depth estimator for subject-agnostic scaling. The modular design enables robust and scalable reconstruction of both human and non-human subjects across long, multi-view sequences.
@@ 3. Method overview
To obtain metric-scale reconstructions, we propose two strategies. The first leverages a human body prior, using the SMPL model when people are present in the scene, similar in concept to HSfM (Müller et al., 2025). The second is subject-agnostic, employing a monocular metric depth estimator to recover scale even in scenes without humans (see Sec. 7).
4. Temporal Synchronization
@@
Rather than relying on raw spectrogram cross-correlation, which is computationally demanding for long recordings (Hasler et al., 2009; Shrstha et al., 2007), we use compact audio descriptors to improve efficiency. Specifically, we represent each audio track using Mel-Frequency Cepstral Coefficients (MFCCs), a perceptually motivated and widely adopted feature in audio analysis. MFCCs capture the spectral envelope of the signal and are robust to noise. The reduced set of features allows for efficient and accurate cross-correlation, making them a suitable choice for estimating temporal lag in multi-camera settings.
The accuracy of this method depends on two main factors: the temporal resolution of the MFCC representation and the physical distance between the sound source and each camera. The former is determined by the hop length $H$ used during the Short-Time Fourier Transform (STFT), and the latter by the time delay resulting from different distances to the sound source. Following prior observations (Hasler et al., 2009), we formalize these two sources of error as $\Delta t_{1}$ and $\Delta t_{2}$, respectively. We derive conditions under which subframe synchronization is achievable. Let $f_{s}$ be the audio sample rate, $f_{r}$ the video frame rate, $H$ the hop length, $d_{1}$ and $d_{2}$ the distance from each camera to the sound source respectively, and $c~\approx 343\,\text{m/s}$ the speed of sound in air. Then:
$\Delta t_{1}=\frac{H}{2f_{s}}~\quad~\text{and}~\quad~\Delta t_{2}=~\frac{|d_{1}-d_{2}|}{c}$
@@
$\Delta t_{1}+\Delta t_{2}<~\frac{1}{f_{r}}$
@@
$H<~\frac{2f_{s}}{f_{r}}-\frac{2f_{s}|d_{1}-d_{2}|}{c}\quad\text{and}\quad|d_{1}-d_{2}|<c(\frac{1}{f_{r}}-\frac{H}{2f_{s}})$
@@
In typical multi-view camera arrangements (e.g., circular rigs), microphone distances to the sound source vary by only a few meters. For instance, with $|d_{1}-d_{2}|<5$ m, $f_{s}=48$ kHz, and $f_{r}=60$ Hz, the hop length $H$ must be below 200 samples to ensure subframe accuracy. Inversely, with $H=1$, the method remains accurate for inter-microphone distances below approximately 5.7 m. In case finer alignment is required, image-based synchronization techniques (Vo et al., 2022; Shuai et al., 2022) can be used to refine an initial audio-based estimate. In all experiments presented in this work, the camera setups fall within our constraints, and audio-based synchronization is sufficient to achieve subframe accuracy. As a result, we did not incorporate image-based refinement in our pipeline.
5. Graph-based Automatic Camera Calibration
@@
This graph formulation enables global reasoning about pairwise relationships and allows us to select an optimal subset of calibrations through a minimum spanning tree (MST). This contrasts with prior incremental approaches that add cameras greedily to a global frame, at the risk of yielding a suboptimal global configuration (Xu and Kitani, 2022; Puwein et al., 2015). By composing relative poses along the MST, we recover global camera extrinsics efficiently with a guarantee of minimizing the total calibration error (Section 5.5). The recovered cameras are expressed in arbitrary units, and metric scaling to real-world dimensions is applied in Sec. 7.
5.1. Intrinsics Initialization
@@
To construct the camera graph, we need to recover the relative poses between camera pairs, which requires decomposing the essential matrices into their corresponding rotations and translation directions. Computing the essential matrix, however, depends on knowledge of the camera intrinsics $K$. In our scenario, where cameras are assumed to be static and the subject is moving, Kruppa’s formulation (Kruppa, 1913; Hartley, 1997) cannot be applied. One alternative is to compute fundamental matrices, which do not require intrinsics, and then optimize the intrinsics for each camera by minimizing the Sampson distance over all edges. However, full-graph optimization is computationally expensive, growing exponentially with the number of cameras. To maintain a linear cost with respect to the number of cameras, we instead use MoGe (Wang et al., 2025b) to obtain an initial estimate of each camera’s intrinsics. This model is later reused to predict scene point maps. These intrinsics are later refined during bundle adjustment (Sec. 5.6), to which we incorporate lens distortion coefficients that follow the Brown-Conrady model, a step omitted in prior works.
5.2. Efficient Pairwise Calibration
@@
We propose a novel approach for pairwise camera calibration that is both robust and scalable to long sequences due to its fixed computational cost. Our method builds upon the work of Puwein et al. (Puwein et al., 2015) by introducing confidence-driven spatio-temporal subsampling of keypoint correspondences, which enhances robustness while maintaining efficiency. Additionally, we leverage a graph-based formulation to compute relative scales reliably, incorporating the quality of each pairwise calibration to produce geometrically consistent relative extrinsics.
@@
The edges of the graph represent pairwise camera calibrations, encoded by an essential matrix $E_{ij}$ that captures the relative rotation $R_{ij}$ and translation direction $\hat{\mathbf{t}}_{ij}$ between cameras $C_{i}$ and $C_{j}$. Classical approaches estimate this matrix using the Structure-from-Motion (SfM) five-point algorithm (Nister, 2004) within a RANSAC loop applied either to all raw keypoint correspondences (Xu and Kitani, 2022) or to a subset of correspondences (Puwein et al., 2015). Both strategies can be unstable when low-quality correspondences are present, and processing all correspondences becomes computationally prohibitive for long sequences.
@@
Learning-based methods such as DUSt3r (Wang et al., 2024; Müller et al., 2025) predict relative poses in a single-frame multiview setting and can be more robust when high-confidence keypoints are unavailable. However, they do not exploit temporal consistency, which can limit accuracy, and their reliance on an additional neural network increases computational overhead; for example, DUSt3r requires approximately 40 ms for inference on a single image pair using an H100 GPU. In contrast, our method is more computationally frugal because it fully exploits the same 2D keypoints that are already required for triangulating the final skeleton output. By selectively subsampling these keypoints based on their detection confidence, we improve the stability and accuracy of previous approaches while preserving the efficiency of classical SfM and ensuring scalability to long sequences.
5.2.1. Confidence-Aware Spatio-Temporal Subsampling
@@
Figure 2. Example confidence-driven keypoint subsampling for a pair of views from Human3.6M. Colors indicate the pair confidence score $w_{kl}$, solid lines represent selected correspondences, and dashed lines represent correspondences that were not selected. The sampling favors keypoints for which the 2D keypoint detector is highly confident about their location (e.g., unoccluded and well-localized keypoints).
@@
An effective spatio-temporal subsampling strategy for keypoint correspondences is important for three reasons: (1) it improves the robustness of essential matrix estimation by filtering out low-quality correspondences and by leveraging information across multiple frames, (2) it reduces computational complexity by limiting the number of correspondences, which is particularly important for scalability on long sequences, and (3) it preserves sufficient geometric diversity to ensure accurate essential matrix estimation. Traditional RANSAC samples correspondences uniformly, ignoring the varying reliability of detections. As a result, if low-confidence correspondences constitute a majority, they can disproportionately influence hypothesis generation, leading to inaccurate relative pose estimates and requiring more iterations to converge on a valid model. Puwein et al. (Puwein et al., 2015) addressed this by sampling correspondences with higher probabilities for temporally distant correspondences and for those involving different keypoints. However, their strategy does not leverage keypoint confidence scores, which can lead to the inclusion of unreliable correspondences.
@@
Building on their method, we improve the spatio-temporal subsampling strategy by incorporating the 2D keypoints detector confidence scores $w_{i,k}~\in[0,1]$, with $i$ the view index and $k$ the keypoints index, into the selection process. Each point correspondence between camera $C_{i}$ and $C_{j}$ is denoted as $(\mathbf{x}_{i,k},\mathbf{x}_{j,k})$. Its score is computed as the geometric mean of the detector confidences:
(1) $w_{ij,k}=\sqrt{w_{i,k}~\cdot w_{j,k}}$
@@
5.3. Graph Edge Reliability Estimation
@@
$\hat{\mathbf{x}}_{i,k}=K_{i}^{-1}\mathbf{x}_{i,k}~\quad~\text{and}~\quad~\hat{\mathbf{x}}_{j,k}=K_{j}^{-1}\mathbf{x}_{j,k}$
@@
(2) $s_{ij}=\frac{1}{M}\sum_{m=1}^{M}\mathcal{S}(E_{ij},\hat{\mathbf{x}}_{i,m},\hat{\mathbf{x}}_{j,m})$
@@
5.4. Solving relative scales
@@
$R_{j}=R_{i}R_{ij}~\quad~\text{and}~\quad~\mathbf{t_{j}}=R_{i}\cdot(\lambda_{ij}\cdot~\mathbf{\hat{t_{ij}}})+\mathbf{t_{i}}$
In previous work, finding this relative scale $\lambda_{ij}$ is addressed by aligning 3D structures in a common space. Xu et al. (Xu and Kitani, 2022) rescale and align the triangulated bones from each camera pair into a shared coordinate system. In their method, the quality of the pairwise triangulation directly affects the result when computing the relative scales. Methods like DUSt3r (Wang et al., 2024) incorporate the scaling factors into a global alignment loop. The predicted pairwise point maps are then scaled and aligned into a globally consistent 3D reconstruction.
Instead, we solve for the relative scales $\lambda_{ij}$ by exploiting loop closure constraints (Arrigoni et al., 2015). The key idea is that if we follow the relative transformations along a closed cycle of cameras, for example moving from $C_{i}$ to $C_{j}$ to $C_{k}$ and back to $C_{i}$, we should return to the starting point. This concept is illustrated in Fig. 3. Formally, each triplet of cameras yields one linear constraint of the form:
(3) $\lambda_{ij}R_{ki}R_{jk}~\mathbf{\hat{t_{ij}}}+~\lambda_{jk}R_{ki}~\mathbf{\hat{t_{jk}}}+~\lambda_{ki}~\mathbf{\hat{t_{ki}}}=~\mathbf{0}$
We collect all $L$ loop constraints into a homogeneous linear system $A\boldsymbol{\lambda}=\mathbf{0}$, where $A\in\mathbb{R}^{3L\times M}$ contains the coefficients from the constraints, and $\boldsymbol{\lambda}~\in~\mathbb{R}^{M}$ represents the unknown relative scale factors.
@@
(5) $WA\boldsymbol{\lambda}=\mathbf{0}$
@@
(6) $\min_{\boldsymbol{\lambda}}\|W^{\frac{1}{2}}A\boldsymbol{\lambda}\|_{2}^{2}+\mu\|\boldsymbol{\lambda}-\mathbf{1}\|_{2}^{2}$
@@
Figure 3. Example of camera loop closure with incorrect (left) and correct (right) relative scale factors. When the relative scales $\lambda_{ij}$ are incorrect, traversing the cycle $C_{i}\!\to\!C_{j}\!\to\!C_{k}\!\to\!C_{i}^{\prime}$ does not return to the original camera position, resulting in $C_{i}\neq C_{i}^{\prime}$. In contrast, when the correct scales are used, the composition of the relative transformations yields the identity, and the loop closes consistently with $C_{i}=C_{i}^{\prime}$.
5.5. Optimal Absolute Extrinsics